# Synthetic Data Generation in Cybersecurity: A Comparative Analysis

### v1

@@ II Related Work
II-A Synthetic Tabular Data Generation Methods
Synthetic data is an artificially produced dataset that mimics the statistical features of real-world data without disclosing any sensitive information [3]. This data type spans various forms, such as tabular data, images, and text.
Tabular data, organized in a table format with rows and columns, is widely used for ML and DL modeling. It is also crucial for cybersecurity applications due to its structured nature. It is ideal for representing logs, transaction records, and other important information for detecting and mitigating security threats [16]. For example, cybersecurity datasets often contain tabular data in IDS, fraud detection, and anomaly detection, enabling precise analysis and decision-making [33, 34]. Additionally, tabular data’s clear and organized format allows for effectively applying ML and DL algorithms to enhance security measures [22].
The significance of synthetic data in cybersecurity stems from the challenges associated with accessing high-quality, real-world data in this field. Privacy concerns, regulatory limits, and the sensitive nature of cybersecurity information often restrict the availability of data [35]. By facilitating the generation of realistic datasets free of genuine, sensitive information, synthetic data offers a solution to these challenges [36]. This advancement is critical for pushing forward research, enhancing security models, and executing thorough threat analyses.
II-B Applications in Cybersecurity
Synthetic data finds multiple applications in cybersecurity, improving different facets of security management and threat response:
  1. 1.
    Intrusion Detection Systems (IDS): By supplying varied and representative datasets encompassing a range of attack patterns and behaviors, synthetic data aids in the training and testing of IDS, enhancing their detection accuracy and resilience [37]. Research has shown that employing synthetic data techniques on datasets like UNSW-NB15 and CICIDS-17 has significantly improved the performance of deep learning models used in IDS [38].
  2. 2.
    Anomaly Detection: Synthetic data plays a role in simulating typical and atypical behaviors within networks, thereby facilitating the training of systems to detect deviations indicative of potential security incidents [39]. Synthetic data has been used to preserve privacy and detect anomalies in the KDD dataset. Its effectiveness, measured in supervised, semi-supervised, and unsupervised anomaly detection, showed an accuracy of 99.69%, similar to models trained on original data [40].
  3. 3.
    Security Training and Simulation: The value of synthetic data extends to crafting realistic scenarios for security training exercises and simulations. This enables security teams to hone their responses to varied cyber threats without compromising real data [41]. A deep Convolutional Generative Adversarial Network (DCGAN) model was used to generate synthetic cyberattack data and realistic attack simulations to train deep learning models for better detection. The model with those simulated attacks achieved an accuracy of 99.69% on the KDD dataset and 97.93% on the CICIDS-17 dataset [42].
In discussing the significance and uses of synthetic tabular data in cybersecurity, we will focus on the methods used to generate such data effectively. The later sections will delve into the specifics and their architecture. They will also provide an experimental analysis of the prominent standard (non-AI) and AI-based methods for creating synthetic tabular data for cybersecurity IDS.
II-C Metrics for Evaluation
  1. 1.
    Fidelity: Fidelity measures how well the synthetic data replicates the statistical properties of the original real-world data [43]. Therefore, the single most significant metric for evaluating synthetic tabular data generally involves assessing the data’s underlying distribution [44]. This approach is based on the definition of data synthesis to capture the probability distribution function of the original data, thereby creating synthetic data that mirrors these characteristics [45]. However, the evaluation process can be subjective, based on the data production’s objective. As a result, researchers employ various methods to assess synthetic tabular data, considering its structure markedly differs from that of images.
    Unlike synthetic tabular data, the validation of synthetic image-based output is often visual. If the model yields realistic images, it’s deemed successful. Images comprise spatial information, shape, texture, objects, and color [46]. Nonetheless, synthetic data is confined to images and encompasses tabular data [47], characterized by its complexity, multiple interconnected features, hidden patterns, and context-dependent information [48]. Hence, evaluating synthetic tabular data demands a more structured approach.
  2. 2.
    Utility: Utility measures the usefulness of synthetic data for the same tasks or analyses as the original data [49]. To evaluate the quality of the synthetic data, key performance metrics such as accuracy, F1 score, precision, and recall were compared between models trained and tested on real data versus those involving synthetic data[50, 15, 51]. Specifically, the following scenarios were used to rank performance:
    TRTR (Train and Test on Real Data): Baseline performance using real data. TSTR (Train on Synthetic, Test on Real): Measures how well models trained on synthetic data generalize to real data. TRTS (Train on Real, Test on Synthetic): Assesses the consistency of synthetic data in replicating real data behavior. TSTS (Train and Test on Synthetic Data): Evaluates synthetic data’s internal consistency and utility within a self-contained model [52, 53]. By comparing these four scenarios, the utility of the synthetic data can be ranked, providing insight into how well it replicates the practical use of real data [54, 55, 56].
  3. 3.
    Universal metrices: Recent research has examined synthetic tabular data via diverse statistical and ML methods [57, 58, 59]. Nevertheless, these approaches could be more precise regarding suitable evaluation measures. Hence, there’s been a move to define universal metrics to boost universality and interpretability. One such metric is TabSynDex, offering a score between 0 and 1 to gauge the synthetic data’s closeness to real data [60]. However, this metric still needs to improve in selecting the proper evaluation measures. For example, there needs confirmation that TabSynDex’s five main components adequately represent the synthetic data’s completeness or facilitate more accessible validation of generated tabular data. A recent SCH Yang et al. proposal introduces a structured evaluation methodology focused on comparing the distributions of synthetic and real data [44]. This study is pivotal as it aligns with generative AI’s core principle: replicating the probability distribution of the original dataset.
Therefore, by considering the recent research in the domain, this study employs prominent evaluation metrics such as fidelity (statistical similarity), utility, and class balance in its evaluation section.
III Synthetic Data Generation Techniques
Figure 1: Methods for generating synthetic tabular data
III-A Standard Methods (non-AI)
  1. 1.
    Random Oversampling (ROS) is a widely used technique to address class imbalance, a common issue in datasets, especially in domains like cybersecurity where certain types of data, such as network intrusions, are rare compared to normal traffic. ROS works by duplicating samples from the minority class to balance the class distribution, thereby preventing ML models from becoming biased toward the majority class [61]. This technique is particularly relevant when the non-complex small dataset lacks sufficient examples of minority classes.
    In synthetic tabular data generation, ROS is often employed to artificially increase the number of minority class samples, helping models better learn the characteristics of these rare events. However, its application in cybersecurity data, which often involves complex relationships between variables, may be limited if used alone, as it does not generate new, diverse data points [62].
    The basic procedure of ROS is straightforward:
    1. (a)
      Identify the Minority Class: Determine which class has fewer samples.
    2. (b)
      Duplicate Samples: Randomly select samples from the minority class and duplicate them until the class distribution is balanced.
    3. (c)
      Augmented Dataset: The result is a dataset with a balanced class distribution, where the minority class is equally represented as the majority class
    This process does not alter the duplicated samples, meaning the new data points are exact copies of the original samples.
  2. 2.
    Synthetic Minority Oversampling Technique (SMOTE) is an advanced over-sampling method designed to address the limitations of ROS by generating synthetic examples rather than merely duplicating existing ones. Introduced by Chawla et al. in 2002, SMOTE aims to balance the class distribution in datasets by creating new instances of the minority class, which helps improve the performance of ML models, particularly in scenarios where the class imbalance is significant [63]. Unlike traditional over-sampling methods, SMOTE synthesizes new samples by interpolating between existing minority class samples, leading to more generalized decision boundaries for classifiers [64].
    SMOTE is particularly useful in domains such as cybersecurity, where the occurrence of attacks (minority class) is much rarer than normal behavior (majority class). The technique is well-suited for situations where the minority class instances are spread sparsely across the feature space, and duplicating existing samples (as done in ROS) could lead to overfitting [65, 66].
    The SMOTE algorithm operates through the following steps:
    1. (a)
      Minority Class: For each minority class instance, there is a need to select k-nearest neighbors, e.g., k = 5
    2. (b)
      Synthetic Samples: One of the k-nearest neighbors is randomly chosen to generate a synthetic data point. The synthetic data point is created by interpolating between the feature vector of the original data point and that of the selected neighbor. This interpolation involves calculating the difference between the feature vector of the selected neighbor and the original data point. Then, this difference is multiplied by a random number between 0 and 1, and the resulting value is added to the feature vector of the original data point.
    3. (c)
      Augmented Dataset: The result is a dataset augmented with synthetic samples that exhibit variations within the minority class, thereby mitigating the risk of overfitting and improving the model’s ability to generalize.
  3. 3.
    Adaptive Synthetic Sampling (ADASYN) is an oversampling method similar to ROS and SMOTE. However, it introduces a more adaptive approach to generating synthetic data. It was introduced by He et al. in 2008 and focuses on the harder-to-learn examples from the minority class. ADASYN improves on SMOTE by adapting the sampling procedure based on the learning difficulty of each minority class sample. Data points that are harder to classify receive more synthetic neighbors; the rest receive fewer neighbors. [67].
    ADASYN is particularly useful in domains such as cybersecurity, where the rarity of specific attack types (minority class) can cause models to perform poorly [68]. By adaptively generating synthetic samples based on the distribution of the minority class, ADASYN helps models learn the decision boundary more effectively, leading to better generalization and improved detection of minority class instances [69].
    ADASYN operates through the following steps:
    1. (a)
      Identify Difficult Minority Class: ADASYN computes the local density of each minority class by analyzing how many of its nearest neighbors belong to the majority class. Data points surrounded by majority-class neighbors are considered harder to classify.
    2. (b)
      Weight Assignment: Weights are assigned to each minority instance/data point based on difficulty level. Harder-to-classify points receive a higher weight, implying that more synthetic samples will be generated.
    3. (c)
      Synthetic Samples: For each minority instance, ADASYN interpolates between the instances and one of its nearest minority class neighbors. Synthetic samples are created along the line connecting the two instances.
    4. (d)
      Data Balancing: ADASYN adds the generated samples to the original ones, focusing on the problematic region to reduce the class imbalance and improve the model training and learning.
  4. 4.
    Cluster Centroids is an under-sampling technique designed to address class imbalance by reducing the number of majority class instances. Unlike traditional under-sampling methods, which randomly remove majority class samples, Cluster Centroids utilizes a clustering algorithm to identify representative centroids within the majority class. The technique was popularized to preserve the majority class’s distribution and structure while reducing its size, thereby balancing the dataset without losing critical information [70, 71]. This approach is beneficial in scenarios where the majority class is substantial, and random under-sampling could significantly lose valuable data.
    The concept of cluster centroids is applicable in various domains, including cybersecurity. In highly imbalanced datasets, where normal traffic significantly outweighs the attack data, cluster centroids help by grouping similar attack (minority class) samples into clusters and generating synthetic data points at the centroid of each cluster. This helps represent the attack patterns more effectively and exposes ML models to diverse attacks for robust anomaly detection [72].
    Cluster Centroids operates through the following steps:
    1. (a)
      Minority Class Clustering: Minority class data is grouped into clusters using a clustering algorithm, e.g., K-Means clustering. Each cluster contains similar data points from the minority class.
    2. (b)
      Centroid Calculation: For each cluster, the algorithm computes the centroids, the central points representing the average position of the data points within a particular cluster in the feature space.
    3. (c)
      Synthetic Data: Synthetic samples are generated at or around the center of each minority cluster. This ensures that the generated data is representative of the overall distribution of the minority class.
    4. (d)
      Balanced Dataset: Synthetic data is added to the original data to balance class distributions, improving the AI model’s ability to learn from minority classes for higher efficiency.
    Figure 2: Bayesian Network illustrating some key feature dependencies in the NSL-KDD dataset
  5. 5.
    Gaussian Mixture Model (GMM) is a probabilistic model used to represent the presence of sub-populations within an overall population without requiring that an instance belong exclusively to a single sub-population. GMMs are particularly useful for modeling data that exhibit multiple underlying distributions, assuming that all data points are generated from a mixture of several Gaussian distributions with unknown parameters [73]. Each Gaussian component within a GMM represents a cluster characterized by its mean, variance, and mixing coefficient.
    In the context of synthetic tabular data generation, GMMs approximate the data distribution, making them highly applicable for generating new samples that preserve the statistical properties of the original dataset. This is especially useful in cybersecurity, where modeling the intricate patterns of network traffic or attack vectors requires flexible and accurate representation of complex, multimodal distributions [74]. GMMs provide a robust method for capturing these patterns, allowing for generating synthetic data that closely mimic real-world scenarios [75].
    GMM operates through the following steps:
    1. (a)
      Fitting GMM: GMM fits the original data by assuming it comes from the mixture of multiple Gaussian distributions. It estimates these Gaussian components’ parameters (mean, covariance, and weights) using the Expectation-Maximization (EM) algorithm.
    2. (b)
      Estimate Gaussian Parameters: The model assigns data points to different Gaussian components based on their likelihood and calculates the mean (center), covariance (spread), and mixture weight (probability and skewness) for each Gaussian distribution.
    3. (c)
      Sampling: After training the model, synthetic data is generated by randomly selecting a Gaussian component (based on its weight) and then sampling a point from the selected Gaussian using its parametric values.
    4. (d)
      Data Generation: Repeating the sampling process multiple times generates synthetic data that mimics the structure and underlying distribution of the original data. Thus, the generated data points from all components are combined to form the final synthetic dataset. [76].
Figure 3: Variational Autoencoder Structure
    III-B AI based Methods
  1. 1.
    Baysian Network (BN) are probabilistic graphical models that represent a set of variables and their conditional dependencies via a directed acyclic graph (DAG) [77]. Bayesian Networks are particularly effective for modeling complex relationships in data where uncertainty and causality play critical roles [78, 79]. In figure 2, this BN illustrates that target is a central feature, influenced by several other variables, including service_http, service_private, flag_SF, logged_in, and flad_S0. This suggests that target plays a crucial role in determining the system’s behavior, as it is obvious by its binary data type of normal or attack. The BN also shows key conditional dependencies, such as flag_SF being dependent on logged_in and service_domain_u, while dst_host_serror_rate influences flad_S0. Some features, like service_private and service_http, directly impact target but do not further influence other variables. Overall, this network emphasizes the significant joint dependencies between these system features, highlighting target as a critical node and capturing the complex relationships between the various network-related parameters.
    In synthetic data generation, BNs generate data that preserves the inherent dependencies and conditional probabilities observed in the original dataset. This is particularly valuable in cybersecurity, where understanding the intricate dependencies between various network events, user behaviors, or attack patterns is essential for generating realistic synthetic datasets. By capturing these relationships, BNs allow for the generation of synthetic data that not only mimics the distribution of the original data but also maintains the underlying causal structures, making it highly applicable for scenarios requiring interpretable and robust synthetic data [14].
    The process of generating synthetic data using Bayesian Networks involves the following steps [80]:
    1. (a)
      Structure Learning: The first step is to determine the structure of the Bayesian Network, which involves identifying the dependencies between variables. This can be done using expert knowledge or algorithms that learn the structure from data, such as hill climbing or score-based methods [81]. Each node in the network corresponds to a variable, and directed edges between nodes represent conditional dependencies.
    2. (b)
      Parameter Learning: After defining the structure, the next step is to estimate the conditional probability distributions (CPDs) for each variable in the network, given its parent node. This joint probability information is learned from the original dataset.
    3. (c)
      Sampling: After learning the structure and parameters (CPDs), synthetic data is generated by sampling from the Bayesian network. First, the root node is sampled from its marginal distribution, and then the child nodes are sampled based on their conditional probabilities. This process continues throughout the network until all variables are sampled.
    4. (d)
      Data Generation: The sampling process is repeated for the required number of synthetic data points. Each sample is a synthesized record for the dataset. This process ensures that the generated data depicts the dependencies learned from the original dataset.
  2. 2.
    Tabular Variational Autoencoder (TVAE) is a generative model designed to handle the complexities of tabular data, often including a mix of continuous and categorical variables. TVAE has the base idea of autoencoders that consists of two parts: an encoder network that learns to compress high-dimensional data into a low-dimensional, latent spacial representation (the code) and a decoder network that decompresses the compressed representation into the original domain as depicted in Figure 3 [82]. Thus, it extends the traditional Variational Autoencoder (VAE) framework by adapting it to the specific needs of tabular data, ensuring that the generated synthetic data accurately reflects the structure and distribution of the original dataset [83].
    Mathematical process of TVAE for Synthetic data generation:
    1. (a)
      Encoder-Decoder Architecture: TVAE uses an encoder-decoder structure where the encoder $q_{\phi}(z|x)$ maps the input data $x$ to a latent space $z$, and the decoder $p_{\theta}(x|z)$ reconstructs the data from this latent representation. The encoder outputs the parameters of a Gaussian distribution in the latent space:
      $q_{\phi}(z|x)=\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ (1)
      Where $\mathcal{N}(z;\mu_{\phi}(x),\sigma_{\phi}(x)^{2}I)$ represents a multivariate normal (Gaussian) distribution where the latent variable $z$ is assumed to follow a Gaussian distribution with mean $\mu_{\phi}(x)$ and covariance matrix $\sigma_{\phi}(x)^{2}I$. The two parameters, $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$, are functions of the input $x$ and are learned by the encoder network. More precisely, $\mu_{\phi}(x)$ is the mean vector of the Gaussian distribution, parameterized by the input $x$. This vector represents the expected value of the latent variable $z$ for a given input $x$, determining the center of the distribution, and $\sigma_{\phi}(x)^{2}$ is the variance of the latent variable $z$, also parameterized by the input $x$. This variance determines how much the latent variable $z$ can deviate from its mean $\mu_{\phi}(x)$, controlling the spread of the distribution. Specifically, $\sigma_{\phi}(x)^{2}$ is the element-wise square of the standard deviation $\sigma_{\phi}(x)$. Plus, $I$ is the identity matrix used to represent that the latent variables are assumed to be independent, meaning there is no covariance between different latent dimensions. The use of $I$ ensures that the covariance matrix is diagonal, implying that each latent variable has its independent variance, controlled by $\sigma_{\phi}(x)^{2}$.
    2. (b)
      Variational Objective: The objective of training TVAE is to maximize the Evidence Lower Bound (ELBO):
      $\begin{split}\mathcal{L}(\theta,\phi;x)=&\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\\ &D_{KL}(q_{\phi}(z|x)||p(z))\end{split}$ (2)
      Here, $\mathcal{L}(\theta,\phi;x)$ is the objective function (ELBO) to maximize during the training of TVAE. It measures how well the model reconstructs data and regularizes the latent space, and $\log p_{\theta}(x|z)$ represents the reconstruction loss, ensuring that the generated data $\hat{x}$ resembles the input data $x$. $D_{KL}$ is the Kullback-Leibler divergence that regularizes the latent space to follow a prior distribution $p(z)$, typically a standard normal distribution $\mathcal{N}(0,I)$. Plus, $\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]$ is the expected log-likelihood term, representing the reconstruction loss. It ensures that the generated data $\hat{x}$ resembles the input data $x$.
    3. (c)
      Latent Variable Modeling: The latent space $z$ is sampled from the Gaussian distribution parameterized by $\mu_{\phi}(x)$ and $\sigma_{\phi}(x)$. This latent space effectively captures the dependencies and distribution of the original data.
    4. (d)
      Data Generation: To generate synthetic data, samples are drawn from the latent space $z$, and these samples are passed through the decoder $p_{\theta}(x|z)$ to produce synthetic data points $\hat{x}$. The decoder ensures that the synthetic data preserves the statistical characteristics of the original tabular dataset.
    5. (e)
      Categorical and Continuous Data Handling: TVAE incorporates specific mechanisms to handle the peculiarities of tabular data. Standard Gaussian distributions are used for continuous variables, while TVAE often employs a Gumbel-Softmax distribution for categorical variables to allow gradient-based optimization.
    Figure 4: Diffusion Model Flow
  3. 3.
    Diffusion Models, particularly Denoising Diffusion Probabilistic Models (DDPM), have emerged as a powerful class of generative models [84, 85]. Originally developed for continuous data types such as images, these models have been adapted to handle the complexities of tabular data, which often includes a mix of categorical and numerical variables. The application of diffusion models to tabular data is exemplified by the Tabular Diffusion Model (TabDDPM), which has shown state-of-the-art performance in generating high-quality synthetic tabular data [86].
    Mathematical process of diffusion models for tabular data generation:
    1. (a)
      Forward and Reverse Diffusion Processes: The diffusion model begins with a forward process that gradually adds Gaussian noise to the data. Given a data point $x_{0}$, the forward process generates a sequence of increasingly noisy versions $x_{1},x_{2},\dots,x_{T}$, where each step adds a small amount of noise (Figure 4):
      $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (3)
      Here, $\beta_{t}$ is a variance schedule that controls the amount of noise added at each step $t$. The term $\sqrt{1-\beta_{t}}$ represents the decay factor, which ensures that the noise doesn’t overwhelm the signal, and $I$ is the identity matrix indicating that the noise is isotropic.
      The reverse process aims to generate synthetic data by denoising a sample starting from pure noise. The reverse distribution $p_{\theta}(x_{t-1}|x_{t})$ is parameterized by a neural network that predicts the mean and variance at each step:
      $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}^{2}I)$ (4)
      Here, $\mu_{\theta}(x_{t},t)$ is the predicted mean and $\sigma_{\theta}^{2}$ is the predicted variance at time step $t$, both parameterized by the neural network $\theta$. The network parameters $\theta$ are optimized to minimize the Kullback-Leibler divergence between the forward and reverse processes, effectively learning to denoise the data step by step.
    2. (b)
      Loss Function: The training objective for diffusion models in tabular data is typically the variational lower bound (VLB), which can be simplified to the mean-squared error (MSE) between the predicted noise and the actual noise:
      $\mathcal{L}(\theta)=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (5)
      Here, $\epsilon_{\theta}(x_{t},t)$ is the predicted noise by the model at time step $t$, where $x_{t}$ represents the noisy data and $\epsilon_{\theta}$ is parameterized by the neural network $\theta$. The term $\epsilon$ represents the noise added to the data during the forward process. The model minimizes the difference between $\epsilon$ and $\epsilon_{\theta}(x_{t},t)$, effectively learning to reverse the noise and reconstruct the data.
      The expectation $\mathbb{E}_{t,x_{0},\epsilon}$ denotes the averaging over all time steps $t$, initial data $x_{0}$, and noise samples $\epsilon$ applied during the forward process.
    3. (c)
      Handling Categorical and Numerical Features: Diffusion models for tabular data, such as TabDDPM, extend the basic framework to handle mixed data types. Categorical variables are typically processed using multinomial diffusion, where the data is represented as one-hot encoded vectors, and the noise is added in a way that respects the unconditional nature of the data:
      $q(x_{t}|x_{t-1})=\text{Cat}(x_{t};(1-\beta_{t})x_{t-1}+\beta_{t}/K)$ (6)
      Here, $\text{Cat}(x_{t};\cdot)$ represents the categorical distribution, which ensures that the synthetic data adheres to the unconditional nature of the variable. The term $\beta_{t}$ is the variance schedule controlling the amount of noise added, while $K$ represents the number of categories. $(1-\beta_{t})x_{t-1}+\beta_{t}/K$ maintains the correct distribution of categories by adjusting the probability mass of each category at every time step $t$.
      Numerical variables are handled using standard Gaussian diffusion processes. The model learns to denoise both categorical and numerical features in a unified framework, ensuring that the generated synthetic data maintains the relationships and distributions of the original data.
    4. (d)
      Synthetic Data Generation: After training, the model can generate synthetic tabular data by sampling from the learned distribution. The reverse process generates a sequence of less noisy samples from a sample of pure noise until the final synthetic data point is produced. This approach allows the model to create realistic tabular data that captures individual feature distributions and the dependencies between features.
    Figure 5: Original GAN structure by Goodfellow [9]
  4. 4.
    Generative Adversarial Networks (GAN)
Conditional Tabular GAN (CTGAN)
CTGAN is a generative model designed explicitly for tabular data, which is often characterized by the presence of mixed data types, including both categorical and continuous variables [12]. CTGAN extends the traditional GAN framework (Figure 5) by incorporating techniques that are well-suited for the unique challenges of tabular data generation, such as handling imbalanced datasets and preserving the relationships between variables [9, 3].
Mathematical process of CTGAN for synthetic data generation:
  1. 1.
    Generator and Discriminator Networks: CTGAN uses a generator network $G(z,c)$ that takes as input a noise vector $z$ sampled from a standard normal distribution and a conditional vector $c$ that encodes the categorical features. The generator produces synthetic data samples that mimic the real data distribution. The discriminator network $D(x,c)$ is then tasked with distinguishing between real data samples $x$ and synthetic samples generated by $G(z,c)$ while also considering the conditional vector $c$.
  2. 2.
    Mode-Specific Normalization for Continuous Data: To handle the diverse distribution of continuous features in tabular data, CTGAN introduces a mode-specific normalization technique. This approach transforms each continuous variable by centering it around a randomly sampled value from its distribution and normalizing it based on the empirical standard deviation of the data within that mode. This process ensures that the generator can more effectively model continuous data distributions:
    $x_{normalized}=\frac{x-x_{sample}}{\sigma(x_{mode})}$ (7)
    where $x$ is the original continuous variable, $x_{sample}$ is a value randomly sampled from the distribution of $x$, and $\sigma(x_{mode})$ is the empirical standard deviation of $x$ within its mode.
  3. 3.
    Conditional Vector Construction: CTGAN utilizes a conditional vector $c$ to guide the generation process, particularly for categorical variables. This vector is one-hot encoded, allowing the generator to focus on producing samples that match the specific conditions defined by $c$. The conditional vector is also integrated into the discriminator’s training process, helping it better differentiate between natural and synthetic data based on the specified conditions.
  4. 4.
    Training Objective: The training process of CTGAN follows the standard GAN framework, where the generator and discriminator are trained in a min-max game. The generator aims to minimize the following loss function to fool the discriminator, while the discriminator tries to maximize its ability to distinguish real from synthetic data:
    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{x\sim p_{data}}[\log D(x,c)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z,c),c))]\end{split}$ (8)
    Where $G$ represents the generator, $D$ represents the discriminator, $x$ is a data sample drawn from the real data distribution $p_{data}$, $z$ is a random noise vector drawn from the prior distribution $p_{z}$, $c$ denotes any additional conditional information, $G(z,c)$ does the generator generate the synthetic data, and $D(x,c)$ and $D(G(z,c),c)$ are the discriminator’s outputs for real and synthetic data, respectively.
  5. 5.
    Data Sampling and Imbalance Handling: CTGAN employs a training-by-sampling technique to address class imbalance in tabular data. The data points are sampled during training to ensure underrepresented classes are adequately represented in the training batches. This approach mitigates the impact of imbalanced datasets on the quality of the generated data.
Copula GAN
CopulaGAN is a generative model that combines the strengths of copula-based statistical models with the flexibility of GANs to generate synthetic tabular data [87]. Copulas are mathematical tools that describe the dependence structure between random variables, enabling the modeling of complex dependencies separately from the marginal distributions [88]. By integrating copulas into the GAN framework, Copula GAN effectively captures the intricate dependencies in tabular data, making it particularly suitable for scenarios where preserving the relationships between variables is crucial [89].
The mathematical process of Copula-GAN for synthetic data generation:
  1. 1.
    Marginal Distribution Estimation: The first step in Copula GAN involves estimating the marginal distributions of each variable in the tabular dataset. This can be achieved using kernel density estimation for continuous variables or empirical distributions for categorical variables. The goal is to transform the original data into a uniform distribution on the interval $[0,1]$ using the cumulative distribution function (CDF) of each variable:
    $u_{i}=F_{i}(x_{i})\quad\text{for }i=1,\dots,d$ (9)
    where $u_{i}$ is the transformed variable, and $F_{i}(x_{i})$ is the CDF of the $i$-th variable.
  2. 2.
    Copula Function for Dependency Modeling: Copula GAN models the dependencies between the variables using a copula function once the data is transformed into the copula space. The copula function $C$ captures the joint distribution of the transformed variables:
    $C(u_{1},\dots,u_{d})=\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ (10)
    where $C$ denotes the copula function, $u_{1},\dots,u_{d}$ are the values of the copula-transformed variables, and $U_{1},\dots,U_{d}$ are the random variables corresponding to these transformed values. The expression $\mathbb{P}(U_{1}\leq u_{1},\dots,U_{d}\leq u_{d})$ represents the joint probability that $U_{1}$ is less than or equal to $u_{1}$, $U_{2}$ is less than or equal to $u_{2}$, and so on up to $U_{d}$ and $u_{d}$.
    The copula function separates the dependency structure from the marginal distributions, which is advantageous for generating synthetic data that preserves the original data’s relationships.
  3. 3.
    Generator and Discriminator Networks: In Copula GAN, the generator $G(z)$ takes a noise vector $z$ and generates samples in the copula space, aiming to mimic the dependency structure captured by the copula function. The discriminator $D(u)$ evaluates whether a given sample from the copula space is real or generated, guiding the generator to produce more realistic samples:
    $\begin{split}\min_{G}\max_{D}&\mathbb{E}_{u\sim p_{data}}[\log D(u)]+\\ &\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z)))]\end{split}$ (11)
  4. 4.
    Inverse Transformation and Synthetic Data Generation: After training, the generator produces samples in the copula space. These samples are then transformed back to the original data space using the inverse of the marginal CDFs:
    $x_{i}=F_{i}^{-1}(u_{i})\quad\text{for }i=1,\dots,d$ (12)
    Where $x_{i}$ represents the $i$-th component of the synthetic data in the original data space, $u_{i}$ is the $i$-th element in the sample in the copula space, and $F_{i}^{-1}$ denotes the inverse of the marginal CDF of the $i$-th variable. This inverse transformation ensures that the generated data points follow the original marginal distributions while maintaining the dependencies modeled by the copula.
Naive Bayes and Logistic Regression GAN (GAN- BLR++
[…]

### v2

@@ II Related Work
The generation of synthetic network traffic data has gained significant attention due to challenges associated with using real-world datasets, such as privacy concerns, ethical limitations, and data imbalance. Existing approaches for synthetic data generation can broadly be classified into Non-AI methods and State-of-the-Art (SOTA) AI methods.
  • Non-AI Methods for Synthetic Data Generation:
    Non-AI approaches primarily focus on statistical techniques and rule-based models for generating synthetic tabular data. Common methods include ROS (Random Over-Sampling), SMOTE (Synthetic Minority Over-sampling Technique), ADASYN (Adaptive Synthetic Sampling), GMM (Gaussian Mixture Models), and Cluster Centroids (CC) [22]. These methods are often applied in scenarios addressing class imbalance rather than comprehensive traffic simulation [36]. For instance, SMOTE and ADASYN have been used to oversample minority classes within network traffic datasets and are frequently compared with modern AI methods such as Variational Autoencoders (VAEs) and GANs (Generative Adversarial Networks) [37].
    Hybrid approaches such as SMOTE-GAN, COPULA-GAN, and Bayesian GAN (GANBLR) [38, 39] have also been explored, combining statistical and generative AI techniques to enhance synthetic data fidelity [40, 41, 42]. Despite their effectiveness in managing class imbalance, these approaches fall short in capturing the complex temporal and structural dependencies of network traffic [21, 43].
  • State-of-the-Art AI Methods:
    AI-based methods dominate the field of synthetic data generation, particularly in recent years. Techniques such as GANs, VAEs, and Diffusion Models have emerged as the leading approaches.
    • GANs (Generative Adversarial Networks): GAN-based models like CTGAN [32], WGAN with gradient penalty [16, 44], CASTGAN [35], GANBLR++ [38, 34] and have been widely used for synthetic tabular data generation. For network traffic-specific datasets, models such as FlowGAN [4], STAN [3], NetShare [6], DoppelGANger [45], and 5GT-GAN [46], MirageNet [5], have shown significant promise.
    • VAEs (Variational Autoencoders): VAEs, particularly TVAE (Tabular VAE), are used as alternatives to GANs for tabular data generation. Recent studies, such as XIDINTFL-VAE, have demonstrated the application of VAEs in handling class imbalance within network traffic datasets [37].
    • Diffusion Models: These models, including TABDDPM [31], NetDiffus [47], and NetDiffusion [7], represent an emerging paradigm in synthetic data generation. The recent NetDiffus framework has outperformed GAN-based methods, providing a 66.4% increase in the fidelity of generated data and an 18.1% increase in downstream ML tasks. They have been reported to outperform GANs and VAEs in terms of data fidelity and diversity.
    • Simulators and Transformer Models: A recent study examined the feasibility of state space models, specifically Mamba, for generating packet-level synthetic network traffic data [48]. They compared the simulator with the state-of-the-art AI models, including a transformer-based TrafficGPT [49].
While these methods have individually shown great potential, they often lack direct comparative analysis under standardized evaluation metrics, making it difficult to identify the most suitable approach for specific network traffic data and its applications. Thus, a significant limitation in the current body of research is the absence of direct comparative studies evaluating Non-AI and AI (conventional and generative/SOTA) methods under unified experimental settings and standardized evaluation scenarios.
From the reviewed literature, the following key research gaps are identified:
  • Limited comparative analysis of non-AI and AI (conventional and generative/SOTA) methods.
  • Lack of standard evaluation metrics to assess synthetic data quality across different approaches.
  • Class imbalance issues remain unexplored in the context of network traffic datasets.
This study aims to address these gaps by systematically comparing non-AI and AI (conventional and generative/SOTA) methods using standardized evaluation metrics of fidelity and utility, with a specific focus on handling class imbalance in synthetic network traffic datasets. Therefore, for the comprehensive comparative study, we have selected twelve prominent synthetic tabular data methods. Technical details of these methods are in the next section.
III Synthetic Network Traffic Data Generation Methods
The generation of synthetic network traffic data can be broadly categorized into Non-AI (statistical) methods and AI-based methods, which include both classical AI and generative AI approaches (Figure 1). This section provides a structured overview of these methods and explains the comparative evaluation approach used in this study.
III-A Overview of Synthetic Data Generation Methods
Non-AI Methods (Statistical Approaches): These methods rely on mathematical models to generate synthetic data by manipulating existing samples or approximating underlying distributions. They are widely used due to their simplicity and low computational cost but often struggle to capture complex traffic dependencies [22, 50].
AI-Based Methods (Classical AI and Generative AI): AI-driven approaches aim to model network traffic more effectively by learning patterns from real data. Classical AI methods use probabilistic models and neural architectures to approximate traffic distributions, while generative AI techniques, such as GANs and diffusion models, focus on generating highly realistic synthetic data through adversarial training or iterative refinements [51, 52].
[…]